the ratio of boys to girls in a class is 4:5 what fractions are boys

Found 2 solutions by greenestamps, Alan3354:
Answer by greenestamps(13334)



Informally....

For every 4 boys, there are 5 girls.
Therefore, out of every 9 students, 4 are boys.
Therefore, 4/9 of the students are boys.

Formally....

The ratio of boys to girls is 4:5.
So let 4x be the number of boys and 5x be the number of girls.
Then 4x+5x = 9x is the total number of students.
Then the fraction of boys in the class is 4x/9x = 4/9.

Answer by Alan3354(69443)

4 times something + 5 times the same something = total B's and G's
4x + 5x = total = 9x
-----
Boys are 4x/9x
= 4/9